Consider three photons of light A,B, and C. Show with calculations the relative order of the energies of the photon in J.
A is blue light of 450 Nm
B is light with frequency 6.46•10^-15 1/s
C is light with energy 3.46•10^-18 J
1
Expert's answer
2020-11-17T10:43:36-0500

"E= hf"

h= 6.63×10-34

A)E=( 6.63×10-34)×(4.5×10-7)

E= 2.98× 10-40 J

B) E=( 6.63×10-34)×(6.64×10-15)

E= 4.4×10-48J

C)3.46•10^-18 J

The relative order of energies is C, A, B